# Infinity-RoPE: Action-Controllable Infinite Video Generation Emerges From Autoregressive Self-Rollout

### v1

@@ 2 Related Work
Bidirectional Video Generation. Recent years have witnessed remarkable progress in video generation, largely driven by the success of denoising diffusion models [ho2020denoising, song2020denoising, lipman2022flow]. These approaches have evolved from early pixel-space diffusion methods to more efficient latent-space formulations [ho2022imagen, singer2022make, blattmann2023align, blattmann2023stable], supported by increasingly powerful architectures ranging from space–time U-Nets [blattmann2023stable, hong2022cogvideo] to diffusion transformers (DiT) [peebles2023scalable, gupta2024photorealistic, veo, hong2022cogvideo, yang2024cogvideox, openai_sora, wan2025wan, kong2024hunyuanvideo]. These large-scale systems typically operate as bidirectional video diffusion models, meaning they leverage information from both past and future frames during the denoising process. This bidirectional conditioning yields coherent and visually consistent results, making such models well suited for offline synthesis and creative video production. However, their reliance on full temporal context prevents deployment in real-time or streaming scenarios, where future frames are unavailable at inference time.
@@
Autoregressive Video Generation. SkyReels-V2 [chen2025skyreels] integrates diffusion forcing [chen2024diffusion] with reinforcement learning and a non-decreasing noise schedule for infinite-length synthesis, though training is computationally heavy. CausVid [yin2025slow] distilled bidirectional DiT into causal generators using asymmetric DMD [yin2024improved, yin2024one], achieving fast, streaming video synthesis but limited consistency beyond short horizons. Self-Forcing [huang2025self] mitigated this train–test mismatch by conditioning on self-generated frames during training, improving temporal stability but still degrading beyond the teacher’s horizon. Self-Forcing++ [cui2025self] extended this with long-horizon distillation on self-rollouts via a rolling KV cache, enabling minute-scale synthesis without long-video supervision. To further suppress drift, Rolling Forcing [liu2025rolling] jointly denoises consecutive frames with progressively increasing noise and maintains a persistent attention-sink cache for global consistency. LongLive [yang2025longlive] complements this with frame-level autoregression and KV re-cache tuning for interactive generation. Recent alternatives explore different causal formulations. NOVA [deng2024autoregressive] performs non-quantized frame-by-frame and set-by-set prediction, removing vector quantization for continuous latent autoregression but limited in long-context scalability. MAGI-1 [teng2025magi] adopts chunk-wise autoregressive denoising with block-causal attention and parallel chunk generation, achieving scalable streaming video generation but at high infrastructure cost.
In contrast to these methods, which primarily extend temporal horizons via new training procedures, distillation strategies, or large-scale infrastructure, $\infty$-RoPE explores what already distilled models can achieve by reparameterizing temporal RoPE and KV caching at inference time, and can be applied in a plug-and-play fashion on top of existing Self-Forcing variants to enable effectively infinite-horizon, controllable video generation.

### v2

@@ 2 Related Work
Bidirectional Video Generation. Recent years have witnessed remarkable progress in video generation, largely driven by the success of denoising diffusion models [ho2020denoising, song2020denoising, lipman2022flow]. These approaches have evolved from early pixel-space diffusion methods to more efficient latent-space formulations [ho2022imagen, singer2022make, blattmann2023align, blattmann2023stable], supported by increasingly powerful architectures ranging from space–time U-Nets [blattmann2023stable, hong2022cogvideo] to diffusion transformers (DiT) [peebles2023scalable, gupta2024photorealistic, veo, hong2022cogvideo, yang2024cogvideox, openai_sora, wan2025wan, kong2024hunyuanvideo]. These large-scale systems typically operate as bidirectional video diffusion models, meaning they leverage information from both past and future frames during the denoising process. This bidirectional conditioning yields coherent and visually consistent results. However, their reliance on full temporal context prevents deployment in real-time scenarios, where future frames are unavailable at inference time.
@@
Autoregressive Video Generation. SkyReels-V2 [chen2025skyreels] integrates Diffusion Forcing [chen2024diffusion] with RL and a non-decreasing noise schedule for infinite-length synthesis. CausVid [yin2025slow] distilled bidirectional DiT into causal generators using asymmetric DMD [yin2024improved, yin2024one], achieving streaming video synthesis but limited consistency beyond short horizons. Self-Forcing [huang2025self] mitigated this train–test mismatch by conditioning on self-generated frames during training, improving temporal stability but still degrading beyond the teacher’s horizon. Self-Forcing++ [cui2025self] extended this with long-horizon distillation on self-rollouts via a rolling KV cache, enabling minute-scale synthesis. Rolling Forcing [liu2025rolling] jointly denoises consecutive frames with progressively increasing noise and maintains a persistent attention-sink cache for global consistency. LongLive [yang2025longlive] complements this with frame-level autoregression and KV re-cache tuning for interactive generation. Stable Video Infinity (SVI) [li2025stable] bridges the train-test gap via error-recycling fine-tuning, which injects self-generated errors during training. BAgger [po2025bagger] constructs corrective trajectories by reversing the model’s own rollouts, enabling the model to learn recovery from drifted states without reliance on bidirectional teachers. Reward Forcing [lu2025reward] enhances motion dynamics through an EMA sink mechanism and Rewarded Distribution Matching Distillation (Re-DMD), which biases the generator toward high-reward motion regions. Block Cascading [bandyopadhyay2025block] introduces a training-free acceleration strategy that parallelizes block-causal inference by allowing future blocks to begin generation conditioned on the partially denoised latents of preceding blocks. NOVA [deng2024autoregressive] performs non-quantized frame-by-frame and set-by-set prediction, removing vector quantization for continuous latent autoregression but limited in long-context scalability. MAGI-1 [teng2025magi] adopts chunk-wise autoregressive denoising with block-causal attention and parallel chunk generation at high infrastructure cost. In contrast to these methods, which primarily extend temporal horizons via new training procedures, distillation strategies, or large-scale infrastructure, $\infty$-RoPE explores what already distilled models can achieve by reparameterizing temporal RoPE and KV caching at inference time, and can be applied in a plug-and-play fashion on top of existing Self-Forcing variants to enable effectively infinite-horizon, controllable video generation.